OpenStudy (anonymous):

solve the differential equation: y''''=0. its 4 times the derivative of y.

OpenStudy (anonymous):

y'''' = 0 means that y''' = constant = c1 y'' = c1*x + c2 y' = c1*x^2/2 + c2*x + c3 y = c1*x^3/(2*3) + c2*x^2/2 + c3*x + c4 To simplify the equation, pack 1/(2*3) into c1 (a constant) and 1/2 into c2 (yet another constant. y = c1*x^3 + c2*x^2 + c3*x + c4

OpenStudy (anonymous):

yes thats the answer. thanks